OpenStudy (anonymous):

Give the 10th term in the geometric sequence. 2, 6, 18, 54, ... A. 39,366 B. 13,122 C. 4,374

OpenStudy (across):

You are on a roll! Keep the sequence going!

OpenStudy (across):

It really is not that hard to see that this sequence is explained by\[a_1=2,\]\[a_{n+1}=3\cdot a_n.\]

sam (.sam.):

a=2 r=3 Tn=ar^(n-1) T10=(2)(3)^(9)=39366

OpenStudy (anonymous):

^To make it in happier language, A1 = the first term, where an+1 is the term proceeding that. For instance, if A1 = 2, then A2 = 3*A1 where A1 is 2. Then A3 = 3*A2, where A2 was found to be 6. Etc. Hopefully that helps!

OpenStudy (across):

Thank you for the explanation, ilikedtea680. :)